OpenStudy (anonymous):

Given the polynomial P(x)=3x^4−x^3−8x^2−4x a. Is 3x + 2 a factor? b. Find the remaining factors if (3x + 2) is a factor

OpenStudy (anonymous):

@sourwing

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

You have two choices: 1. Divide the polynomial by 3x+2 and see if the remainder is 0 2. Factor the polynomial.

OpenStudy (mertsj):

What is your pleasure, my dear?

OpenStudy (anonymous):

divide @Mertsj

OpenStudy (mertsj):

Did you do it?

OpenStudy (anonymous):

Im not sure how to

OpenStudy (mertsj):

I can't show you because the draw function isn't working.

OpenStudy (anonymous):

can you type it?

OpenStudy (mertsj):

First factor out the x

OpenStudy (anonymous):

ok I did @Mertsj

OpenStudy (mertsj):

Did you get: x(3x^3-x^2-8x-4)

OpenStudy (anonymous):

yeah

OpenStudy (mertsj):

So if 3x+2 is a factor then (3x+2)( )=3x^3-x^2-8x-4

OpenStudy (anonymous):

wait what? theres a blank in the middle

OpenStudy (mertsj):

I know. That would be the answer to the long division problem. It would have to start with x^2 and end with -2

OpenStudy (anonymous):

can you show it to me ?

OpenStudy (mertsj):

(3x+2)( x^2 -2 )=3x^3-x^2-8x-4

OpenStudy (mertsj):

I can try.

OpenStudy (anonymous):

ok

OpenStudy (mertsj):

___x^2_-x___-2_________________________ 3x+2| 3x^3-x^2-8x-4 3x^3+2x^2 -3x^2-8x -3x^2-2x -6x-4 -6x-4

OpenStudy (anonymous):

why are there lines?

OpenStudy (anonymous):

ok nvm I know why

OpenStudy (anonymous):

ok so what do I do now?

OpenStudy (mertsj):

Because it is a long division problem

OpenStudy (mertsj):

Answer the questions. Is 3x+2 a factor?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its no right?

OpenStudy (mertsj):

Why do you say no?

OpenStudy (anonymous):

because its not in the answer at the top in long division ?

OpenStudy (mertsj):

What is the answer to the long division problem?

OpenStudy (anonymous):

x^2-x-2

OpenStudy (mertsj):

What is the remainder?

OpenStudy (anonymous):

0

OpenStudy (mertsj):

Is 4 a factor of 12?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

Yes it is because 12 /4 is 3 and the remainder is 0

OpenStudy (mertsj):

So use that same reasoning. Is 3x+2 a factor of your polynomial?

OpenStudy (anonymous):

yes cause the remainder is 0

OpenStudy (anonymous):

right?

OpenStudy (mertsj):

Yes so now we know that your polynomial is equal to : x(3x+2)(x^2-x-2)

OpenStudy (anonymous):

so how can I find the remaining factors for part b ?

OpenStudy (mertsj):

So now do the second part which is to complete the factoring.

OpenStudy (mertsj):

You surely know how to factor a trinomial. x^2-x-2

OpenStudy (anonymous):

factoring of the polynomial right ?

OpenStudy (anonymous):

3x^3-x^2-8x-4

OpenStudy (mertsj):

3x^3-x^2-8x-4=x(3x+2)(x^2-x-2)

OpenStudy (mertsj):

We know those are the factors but the third factor can be factored again. That is your job now...to factor x^2-x-2

OpenStudy (anonymous):

thanks @Mertsj

OpenStudy (mertsj):

did you get it finished?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

Good for you!!

OpenStudy (anonymous):

I got (x+1) (x-2)

OpenStudy (mertsj):

yes So 3x^4--x^3-8x^2-4x=x(3x+2)(x+1)(x-2)

OpenStudy (mertsj):

And that is what you should write on your paper.

OpenStudy (anonymous):

thanks so much!

OpenStudy (mertsj):

yw